OpenStudy (anonymous):

fan and medal...rate of change f(x)=x^4+x^3-5x^2+2x-2 from x=-1 to x=1

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Oh, we have two points: (-1, -9) and (1, -3) We plug these into the average rate of change formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-3+9}{1+1}\) Can you finish that? Just add the top and the bottom, then divide.

OpenStudy (igreen):

@ladystarkzombieslayer

OpenStudy (anonymous):

6/2

OpenStudy (igreen):

Yep, now divide that.

OpenStudy (anonymous):

3/1

OpenStudy (igreen):

Yep, so your answer's 3..or 3/1.

OpenStudy (anonymous):

but the possible answers where 1, -1, 5, -5

OpenStudy (igreen):

Hmm..

OpenStudy (igreen):

I didn't make any mistakes..I don't know what's wrong.

OpenStudy (anonymous):

@Abhisar

OpenStudy (igreen):

@hartnn

OpenStudy (igreen):

What did I do wrong?

OpenStudy (igreen):

https://www.desmos.com/calculator/m5aju3qygr

OpenStudy (igreen):

^^^ That gave me (-1, -9) and (1, -3).

hartnn (hartnn):

i get 3 too ...strange it is not in the choices...

OpenStudy (igreen):

Hmm..maybe you wrote the equation wrong? @ladystarkzombieslayer

OpenStudy (anonymous):

oo im sooo sorry it was to be f(x)= x^4+3x^3-5x^2+2x-2

OpenStudy (igreen):

Okay, that gives us (-1, -11) and (1, -1). Plug them into the formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-1+11}{1+1}\) Simplify that, and what do you get? @ladystarkzombieslayer

OpenStudy (anonymous):

10/2= 5/1=5

OpenStudy (igreen):

Yep, that's your answer.